OpenStudy (anonymous):

The half-life of radium is 1690 years. If 50 grams are present now, how much will be present in 390 years?

OpenStudy (phi):

\[ m_0 2^{-\frac{t}{1690}} \]

OpenStudy (wolf1728):

Ending Amount = Beginning Amount / 2^n where n is elapsed time / half-life Ending Amount = 50 / 2^(390/1690) Ending Amount = 50 / 2^0.2307692308 Ending Amount = 50 / 1.17346046 Ending Amount = 42.6090198214 grams or about 42.61 grams a good web page about half life is http://1728.org/halflife.htm has formulas and a calculator